KyledaGreat:

Consider the following functions. f(x)=3sqrtx and g(x) = sqrtx-1 √ x and g(x)= √ x−1 Find the formula for (f+g)(x) and simplify your answer. Then find the domain for (f+g)(x). Round your answer to two decimal places, if necessar

KyledaGreat:

\[f(x) = \sqrt[3]{x} \] and \[g(x) = \sqrt{x-1}\]

KyledaGreat:

(f+g)(x) = Domain =

KyledaGreat:

@vocaloid

Vocaloid:

it's asking to add the two functions so you can just add f(x) + g(x) to get \[f(x)+g(x)=\sqrt[3]{x}+\sqrt{x-1}\] these are two different types of roots (one's a square root one's a cube root) so you can't really simplify further than this as for domain, the cube root has no limitations, but remember that the square root must be 0 or greater

KyledaGreat:

Domain : [1, ∞)

Vocaloid:

good

KyledaGreat:

i couldn't find the solution at first